OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 19 and the last term is 319,333?

OpenStudy (anonymous):

etermine the three cubic roots of 8.give the answer in the form of a+bj.also give the main root in the exponential form

OpenStudy (anonymous):

dude that sounds so complex...

OpenStudy (anonymous):

Ya cos cubic roots of 8 is 2.2.2 so wanna know how to solve it

OpenStudy (campbell_st):

the 6th term is 319333 you need to find r so \[319333 = 19\times r^{6 -1}\] \[\frac{319333}{19} = r^5\] then\[ r = \sqrt[5]{\frac{319333}{19}}\] Once you know r then use \[S_{6} = \frac{19(r^6 -1)}{r -1}\] that will give you the sum of the 6 terms